by selling a dinner set for rs4320 a crockery dealear made a profit of 8%. find the cp of dinner set?

Grade:7

1 Answers

Arun
6 years ago
Let cost of dinner set = x
Then profit = (4320- x) *100/x
But in question, profit is 8%
Hence
(4320 -x)*100/x = 8
4320*100 - 100x = 8x
108 x = 432000
Hence x = 432000/108
x = 4000 rs
Hence c.p. is 4000 rs